• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Beck

    Log entry time 16:02:47 on August16,2006

    Entry number 180388

    keyword=Shift Summary

    Shift summary
    8:00 No beam due to radiation monitor failure.
    8:45 Start production again continue with kin. point 30.
    Bean current increased to 10uA
    11:00 End of kin. point 30. Trying without success to move
    left arm to 90 deg. according to kin. point 25.
    12:20 Controlled access to repaire/replace left arm bogies encoder.
    16:00 Still hall in controlled access due to the bogie problem